OpenStudy (anonymous):

can anyone help me with this differential eq please t dy/dt + 4y^2=0 t>0

OpenStudy (ribhu):

rearrange the equation separate the variables.

OpenStudy (anonymous):

is it ok if I leave the -4 on the right side

OpenStudy (solomonzelman):

you mean on the right side? sure, wy not...

OpenStudy (anonymous):

After doing that I got -y^-1=-4lnt+c

OpenStudy (anonymous):

So I know the negative would cross out but what about the C would it be -C?

OpenStudy (solomonzelman):

negative would cross out (?), elaborate please...

OpenStudy (anonymous):

uh so I have y= 1/4lnt +C. I'm not sure if the answer is that or y=1/4lnt-C

OpenStudy (solomonzelman):

if you had y^-1 = -4 ln(t) + C then wouldn't y be over that ?

OpenStudy (anonymous):

If i was solving for y for example

OpenStudy (solomonzelman):

y^-1 = 4 ln(t) + C raise both sides to -1 exponent, y = 1 / ( -4 ln(t) + C )

OpenStudy (anonymous):

oh I meant that there was a negative before y^-1

OpenStudy (solomonzelman):

to be fair though, integral 1/t = |t| and you don't need t>0

OpenStudy (solomonzelman):

and I have a typo

OpenStudy (solomonzelman):

-y^-1 = 4 ln(t) + C y^-1 = - 4 ln(t) + C (+C or -C, you are still adding an arbitrary constant) y = 1 / ( - 4 ln(t) + C )

OpenStudy (anonymous):

gotcha thank you so much :)

OpenStudy (solomonzelman):

this is your y function apparently

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/go2zsfuum4

OpenStudy (solomonzelman):

** set of functions